Excel Online Structured Activity: Corporate valuation
Dantzler Corporation is a fast-growing supplier of office products. Analysts project the following free cash flows (FCFs) during the next 3 years, after which FCF is expected to grow at a constant 4% rate. Dantzler's WACC is 10%.

Year   0   1   2   3
.......   .......   .......   .......   .......   .......   .......   .......
FCF ($ millions)   .......   .......   .......   .......   .......   .......   .......   ......
- $14   $15   $42
The data has been collected in the Microsoft Excel Online file below. Open the spreadsheet and perform the required analysis to answer the questions below.

Open spreadsheet
What is Dantzler's horizon, or continuing, value? (Hint: Find the value of all free cash flows beyond Year 3 discounted back to Year 3.) Round your answer to two decimal places. Enter your answer in millions. For example, an answer of $13,550,000 should be entered as 13.55.
$
million

What is the firm's value today? Round your answer to two decimal places. Enter your answer in millions. For example, an answer of $13,550,000 should be entered as 13.55. Do not round your intermediate calculations.
$
million

Suppose Dantzler has $189 million of debt and 29 million shares of stock outstanding. What is your estimate of the current price per share? Round your answer to two decimal places. Write out your answer completely. For example, 0.00025 million should be entered as 250.
$
Answer #1

1.
=42*1.04/(10%-4%)=728.00

2.
=-14/1.1+15/1.1^2+42/1.1^3+42/1.1^3*1.04/(10%-4%)=578.18

3.
=(-14/1.1+15/1.1^2+42/1.1^3+42/1.1^3*1.04/(10%-4%)-189)/29=13.42